Ammonia exists in two forms, ammonium ion (NH4+) and ammonia gas (NH3). In solution, they are related by the reaction Nh4=H+NH3 pK=9.24 If the pH is 8.5 and the total ammonia concentration (ct) is 40 mg/L as N, calculate the concentration of dissolved ammonia gas (NH3) concentration id mg/L (Answer is 7.33 mg/L )
Answer #1

pH = pka + log(dissolved(NH3)/ammonia)

concentration of dissolved NH3 = X/17 mM

concentration of un dissolved NH3 = 40/17 mM


8.5 = 9.24+log((x/17)/((40/17)))

x = concentration in mg/L = 7.3 mg/L